OpenStudy (anonymous):

Use the quadratic formula to find the zeros of this function. f(x) = 2x2 + 6x – 20 Write your answer as two integers separated by the word and, like this: 12 and -3

OpenStudy (anonymous):

recall that the quadratic equation is\[\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]and this is used to find when equations of the form\[ax^2+bx+c=0\]where a,b, and c are all constants so by using your equation, plug in a=2,b=6, and c=-20 into the quadratic equation to get your 2 answers, so you just need to solve this \[\frac{ -6 \pm \sqrt{6^2-4(2)(-20)} }{ 2(2) }\]

OpenStudy (anonymous):

what would that be ?

OpenStudy (anonymous):

\[\frac{ -6 \pm \sqrt{6^2-4(2)(-20)} }{ 2(2) }=\frac{ -6 \pm \sqrt{36+160} }{ 4 }=\frac{ -6 \pm \sqrt{196} }{ 4 }\frac{ -6 \pm 14 }{ 4 }\]therefore your answers come out as\[\frac{ -6+14 }{ 4 }=2\]and\[\frac{ -6-14 }{ 4 }=-5\]